OpenStudy (anonymous):

What are the fourth roots of sqrt 3 + i

OpenStudy (lgbasallote):

i mean \[\LARGE \sqrt[4]{\sqrt 3 + i}\]

OpenStudy (lgbasallote):

is that the question?

OpenStudy (chaise):

I don't think so, lol. It's asking for roots in the complex plain. Humiliated yourself again :)

OpenStudy (anonymous):

@lgbasallote is right,...:)

OpenStudy (anonymous):

it just asks what are the fourth roots of \[\sqrt{3} + i\]

OpenStudy (anonymous):

put it in trigonometric form first

OpenStudy (chaise):

Convert \[\sqrt3+i\] Into -> \[|v|(\cos \theta +i \sin \theta)\] or \[|v|cis \]

OpenStudy (chaise):

Once it is converted, use demvoirs thereorm in order to do solve.

OpenStudy (anonymous):

I've always found \(\text{cis}(x)\) to be such a lame idea.

OpenStudy (anonymous):

\[\sqrt[4]{2cosPi/6+isinPi/6}\]

OpenStudy (chaise):

It's only useful cause you can effectively use demoices therom.

OpenStudy (anonymous):

that's the point

OpenStudy (anonymous):

Define "effectively"

OpenStudy (anonymous):

I have always use deMoivre's theorem without \(\text{cis}(x)\).

OpenStudy (chaise):

Well, go ahead ^_^ @myko z^4=2cos(pi/6)+isin(pi/6) Now use rules of demoivre's theorm.

OpenStudy (unklerhaukus):